[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Archives]

GSFC XMM-Newton GOF Status Report #044: XMM-NEWTON NEWS #13


                 ESA, XMM-Newton Science Operations Centre at
             Villafranca del Castillo, Satellite Tracking Station
                 P.O. Box - Apdo. 50727, 28080 Madrid, Spain


SOC Home Page:           http://xmm.vilspa.esa.es/
Helpdesk web interface:  http://xmm.vilspa.esa.es/xmmhelp
Helpdesk email address:  xmmhelp@xmm.vilspa.esa.es
News Mailing List:       http://xmm.vilspa.esa.es/news/xmm_news_signup.html


           - Reminder: AO2 deadline
           - EPIC-PN status report

Dear Colleagues,

Reminder: AO2 deadline

The deadline for the submission of XMM-Newton AO-2 proposals is getting 
closer. As announced in XMM-Newton News #10, proposal submissions via the
XMM-Newton Remote Proposal Submission system (XRPS) will have to be 
finalized by

25-October-2001 12:00 UT

At the time of writing you have approximately 48 hours to go!!

In view of possible network congestion close to the deadline, proposers 
are strongly advised to submit their proposals sufficiently in advance 
of the due date.

EPIC-PN status report

Why an EPIC-PN status report will you ask yourself?

On 17-Sep-2001 at 15:27 a telemetry parameter of the EPIC-PN
camera went out of limits (the alarm limit is 400 mA and the parameter
reached 478 mA). The nominal value observed for this parameter is around
300 mA. The related CRP (Contingency Recovery Procedure) was promptly
executed and at 15:29 the instrument was commanded to Safe stand-by.
Analysis of the science data produced at the time of the over-current
showed that the camera was still functioning correctly until it was
switched off.
A test was performed on 24-Sep-2001 where only a part of the EPIC-PN
camera was switched on. Following analysis of the acquired housekeeping
data, a further test was executed on 1-Oct-2001 where also the CCD's
were eventually switched on. After some time an internal calibration
exposure was started.
Although all of these tests were successful, it was decided by the
project scientist to await the end of the eclipse season for the full
resumption of EPIC-PN operations, as at that point in time the EPIC-PN
camera can be switched on and left on for approximately 5 months
(until the next eclipse season).
The EPIC-PN camera was switched on after the eclipse season on
14-Oct-2001, and the scientific performance of all quadrants is
as good as it was before.

Investigations into the exact nature of the observed events will
continue, but the EPIC-PN camera is, once again, producing data of
excellent quality!

Some observers will have noted significant changes in the planned
observations over the 17-Sep - 14-Oct timeframe, and you will
now realise this had to do with optimising the observing programme
whilst not using EPIC-PN. In most cases modified observing strategies
were agreed with the PI's of proposals observed in that period, but
should you have any further queries, please contact us at:


With best regards,

The XMM-Newton SOC.

XMM-NEWS mailing list